OpenStudy (crisulcampo):

Is I have a rectangular prism which basis area is 12 cm^2 and height 8 cm, then calculate the values of "a" and "b" where "a" and "b" are basis dimensions... The best medal ever

OpenStudy (crisulcampo):

@pooja195

jhonyy9 (jhonyy9):

a rectangular prism has a base of rectangle area of a rectangle is ??? do you know this formula ?

OpenStudy (crisulcampo):

area= a*b

jhonyy9 (jhonyy9):

exactly so you need getting two integers a and b with product equal 12

jhonyy9 (jhonyy9):

ok. one choice may be 12 = 3*4 right ?

OpenStudy (crisulcampo):

look at that image

OpenStudy (crisulcampo):

I need to find Lateral area, total area and volume

jhonyy9 (jhonyy9):

this image dont change nothing for you get this a and b

OpenStudy (crisulcampo):

I found volume value which is 96 cm^3

jhonyy9 (jhonyy9):

yes sure this was easy bc. V = area of base time height so = 12*8 = 96 cm^3

OpenStudy (crisulcampo):

12= 6*2 12= 4*3 12= 12*1 12= 2sqrt(3)*2sqrt(3)

jhonyy9 (jhonyy9):

2sqrt3 not can being bc. this is a rectangle not a square

OpenStudy (crisulcampo):

I already know the answers, but I just wanna know how to get the answer no making assumptions

jhonyy9 (jhonyy9):

and so the a need being different the b

OpenStudy (crisulcampo):

do you get me?

OpenStudy (crisulcampo):

that's right! a is different than b

jhonyy9 (jhonyy9):

bc. in this text of your exercise given just these details so there are much possibility how you ve wrote above 3*4 6*2 12*1

OpenStudy (crisulcampo):

Could you find a formula that relates basis area with that basis rectangle perimeter?

jhonyy9 (jhonyy9):

why ? do you know the perimeter too ?

jhonyy9 (jhonyy9):

area = a*b perimeter = 2(a+b)

OpenStudy (crisulcampo):

I don't xD... That's why I want to find that formula... I already have the area value

OpenStudy (crisulcampo):

I guess that the only way to solve this exercise is by making assumptions of "a" and "b" values

OpenStudy (crisulcampo):

don't you think?

jhonyy9 (jhonyy9):

ok. a*b = 12 => a = 12/b P = 2a +2b =2*(12/b) +2b do you think it so ?

OpenStudy (crisulcampo):

\[perimeter = \frac{ 24 }{ b }+2b\] \[perimeter = \frac{ 24+2b^2 }{ b }\] \[b*perimeter=24+2b^2\] \[2b^2-b*perimeter+24=0\] \[b=\frac{ perimeter ∓ \sqrt{(-perimeter)^{2}-4*2*24}}{ 4 }\]

OpenStudy (crisulcampo):

I still have an unknow value which is perimeter... @jhonyy9

OpenStudy (sshayer):

your statement is incomplete. dimensions of base may be 1*12,2*6,3*4 if a and b are integers. a*b=12 p=2(a+b) =2(1+12)=26 or =2(2+6)=16 or =2(3+4)=14 if a and b are real numbers ,then there are infinite real numbers (a neq 0,b neq 0) such that ab=96 a=96/b p=2(a+b)=2((96/b)+b)

OpenStudy (sshayer):

in your case b is real so\[P^2-4*2*24 \geq 0\] \[P^2 \geq 4*16*3\] \[\left| P \right|\geq 8\sqrt{3}\] so \[P \leq-8\sqrt{3}~or P \geq 8\sqrt{3}\] perimeter can not be negative Hence \[P \geq 8\sqrt{3}\] so minimum value of perimeter is\[8\sqrt{3}\]

OpenStudy (crisulcampo):

thanks a lot @sshayer :)

OpenStudy (sshayer):

correction write 12 in place of 96

OpenStudy (sshayer):

yw